OpenStudy (anonymous):

7x^2-13y=0 ; solve for X

OpenStudy (saifoo.khan):

u ther?

OpenStudy (anonymous):

yeah

OpenStudy (saifoo.khan):

k, cool.

OpenStudy (saifoo.khan):

7x^2 = 13y Cross multiply x^2 = \[\frac{13y}{7}\]

OpenStudy (anonymous):

well, it says the answer should be: x=±√91y/7

OpenStudy (saifoo.khan):

Put roots to cancel square: FInal Answer: \[\sqrt{\frac{13y}{7}}\]

OpenStudy (saifoo.khan):

it cant be 91.

OpenStudy (anonymous):

hm...I think I'm supposed to do something like let U= blah and U^2= Blah or something...

OpenStudy (saifoo.khan):

umm, nope.

OpenStudy (anonymous):

my notes say take an equation that is not quadratic and write it in to equivilant quadratic form then use quadratic forumla to find solutions. But, I don't think this is one of those problems

OpenStudy (anonymous):

Im quite confused about it >.< but that is the answer my teacher had there? maybe he is wrong!

OpenStudy (saifoo.khan):

lol, your notes are confusing u + they are confusing me too!

OpenStudy (anonymous):

well, it is a test review, so it doesn't say which part of my notes to use. But there is a section of the test lower that asks me to do what I said above, so I don't think this is that. and when I worked it myself, I got what you got. I wonder if he just put the wrong answer?

OpenStudy (saifoo.khan):

Might Be.

OpenStudy (anonymous):

will you help me with this one instead: \[\sqrt{a^2+b^2}=C ;a\]

OpenStudy (anonymous):

solve for a

OpenStudy (anonymous):

it was supposed to space out. it's =c (solve for a)

OpenStudy (saifoo.khan):

Oh, okay! got that. :)

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

\[\sqrt{C^2 - b^2 }\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (saifoo.khan):

\[a = \sqrt{C^2 - b^2 }\]

OpenStudy (saifoo.khan):

well, can u join me on twiddla? i can show u the working there!

OpenStudy (anonymous):

twiddla? Ive never heard of it. lemme see

OpenStudy (saifoo.khan):

http://www.twiddla.com/578717